CLOSURE

TE[[Apply (F, X)]] = \{ \begin{matrix}n & = & P select (F, Arity); \\ fire_{b} & = & Equal? (n, 1); \\ fire_{i} & = & BooltoInt(fire_{b}); \\ res & = & Dispatch_{2,2} (fire_{i}, \\ & \{ \begin{matrix}fun & = & P select (F, Functioname); \\ as & = & P select (F, Chain); \\ as′ & = & Ap_{2} (Arg chain, X, as); \\ res′ & = & Ap (fun, as′); \\ In res′\}, \end{matrix} \\ & Ap_{2} (Make closure, F, X)) \\ In res\} \end{matrix}

TE[[Ap_{n,m} (F, X^{→})]] = {_{m} f′ = P select (F, Fastcallname); t_{m}^{→} = Ap_{n,m} (f′, X^{→}); In t_{m}^{→} }